Sol Summary – February 10th

Sol: Sol #6
Summary Title: The sun came out for our EVA #7
Author’s name: Annalea Beattie
Mission Status: Nominal
Sol Activity Summary: It snowed overnight but dried off quickly around the hab and the sun came out. We all did yoga led by Rajvi. After breakfast we spent the morning discussing the Mid-Mission Report. We waited for Mission Support to confirm our proposed EVA #7 (which was pending, weather permitting). An hour before we were due to leave, we were given the go ahead. Good news for our crew. EVA #7 headed off by foot east of the hab after a quick lunch. This afternoon I compiled the Mid-Mission report and made a fruit crumble. We kept in close communication with our away team. Later Daniel worked in the Science Dome on his samples successfully extracting DNA. More cooking and more work on the Mid-Mission Report followed. We are still busy now with our routine chores, checking the plants, the suits, the systems, all the things we need to do to keep our hab running smoothly.
Look Ahead Plan: Yesterday’s postponed EVA to the south of Cowboy Corner is requested for tomorrow for Clare, Bharti and me with the goals of slope measurements and drawing samples in situ the field.
Anomalies in work: None
Weather: 40F with 58% humidity. Sunny with a High UV for most of the day. Sunset at 5.59pm in the minuses overnight.
Crew Physical Status: Nominal
EVA: EVA #7 to east of the hab to collect water samples for Daniel, concretions, for gypsum for Rajvi and for Clare to monitor different types of weathering.
Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA report, Astronomy Report, EVA request.
Support Requested: None

Operations Report – February 10th

SOL: #6
Name of person filing report: Rajvi Patel
Non-nominal systems: None
Notes on non-nominal systems: Hab Heater turned off
– The Hab heater turned off and had orange (1 short and 3 long) flashes. I contacted Sergii and was able to get the heater back up again.
ROVERS
Spirit rover used: Not used
Opportunity rover used: Not used
Curiosity rover used: Not used
Perseverance rover used: Not used
General notes on rovers:
Summary of Hab operations: put notes here
WATER USE: 24.18 gallons
Water (static tank): 341.09 gallons
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: No
Summary of internet: Nominal
Summary of suits and radios: The radios worked fine for today’s EVA. Suit #6 fan started running extremely low (in the airlock) after returning from the EVA. Our EVA duration was 2.5 hours. The voltage before EVA was 13.12 V and post EVA it was 9.15V. The fan running low might be due to the low voltage. I remember Crew 291 Engineer performed a 1-hour fan run test to see if the battery was faulty for a similar situation. Do I have permission to run that test on Suit #6? Need Advice.
Summary of GreenHab operations: put notes here
WATER USE: 8.6 gallons
Heater: On
Supplemental light: 22:00 -02:00
Harvest: None
Summary of Science Dome operations: Daniel used the Science Dome after sending the EVA team out, starting a non-targeted PCR run of extracted DNA from samples collected on EVA 4 on Sol 4. Towards the end of the run, he prepared an electrophoresis gel to visualize the results of the PCR and found DNA from soil that was collected from underneath a rock. An image was taken and annotated, further PCR’s will be performed to identify this DNA and investigate the presence of a range of biogeochemical genes. Annalea used it for her drawings on concretions.
Dual split: Heat ON
Summary of RAM operations: (Tools used, work done) None
Summary of any observatory issues: None
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support:
– Concern on Suit #6. Refer to the “summary of suits and radios” section.

Journalist Report – February 10th

Sol 6: Outside at last!
Well, it snowed again overnight. And today, none of us were quite so enthusiastic about it, knowing that it would cancel the EVA we had planned for the afternoon. But we kept our hopes up, thinking that if the snow melted and the ground dried enough, we might still be able to go on the EVA.
To ensure we got some exercise either way, we did some yoga in the Hab after a bit of furniture rearranging. Rajvi led the session, and the whole crew joined in. We definitely all felt well stretched and ready for the day ahead afterwards.
The morning was tense, watching and waiting for the snow to melt, all while trying to get some good work done. However, looking out the window every few minutes in anticipation makes it pretty hard to work! By 1130 it was looking fairly dry, and we were feeling hopeful, so at 1200 we radioed Mission Support to check in about the weather. Thankfully, outside was as dry as it looked, and we were given the green light to go ahead with the EVA but proceed with caution as some areas remained muddy.
We hurried to have lunch, and then it was time to put on our EVA suits and go outside! We were looking for concretions that had been found at a site of white mudstone yardangs, which we found after a bit of walking back and forth. The four Mangalyatris on the EVA were enjoying being out and about, as well as actually finding some of our research targets close to the Hab. Our efficiency in the field meant a return before the proposed return time – no frolicking on Mars.
Once back at the Hab the EVA crew discussed the geology and geomorphology they had seen with the resident geologist, Bharti. We also handed over some puddle water to the Crew Biologist Daniel for his continued microbial studies.
While the EVA had been going on, the three back at the Hab certainly hadn’t been idle. Daniel managed to extract DNA from the soil samples returned on Sol 4. What a great success! We’re all patiently awaiting his further PCR tests (after already conducting a non-targeted PCR and gel electrophoresis). Bharti and Annalea helped as HabCom (which Bharti quite enjoyed, given her excitable nature), but equally continued their work on geology and art, respectively. Bharti has a base map coming along nicely, and Annalea’s artworks are uncannily lifelike. The Hab crew also put in some cooking effort to start dinner and dessert preparations (I hear there’s a cobbler on the way).
Sol 6 has felt like it’s ending as soon as it’s begun. All the Mangalyatris are feeling enthusiastic about the possibility of clearer weather and therefore more EVAs! We’re also all excited to see each other’s projects returning results. It’s hard to believe this is already the middle of our mission, but the second half will surely be even more successful than the first half.
Glad to see results coming in,
Clare Fletcher (292 Mangalyatri Journalist)

GreenHab Report – February 10th

GreenHab Officer: Mehnaz Jabeen
Environmental control: heater on , fan off , door close
Average temperature
6:30 70 F, 27%
18:30: 75.F, 20%
Hours of supplemental light: 22:00 – 02:00
Daily water usage for crops: 8 gallons
Daily water usage for research and/or other purposes: 0.6 gallons
Water in Blue Tank (200 gallon capacity): 151.94 gallons
Time(s) of watering for crops: 18:30
Changes to crops: More ripe tomatoes
Narrative:
Yesterday as I mentioned in my report about my new experiment to validate my research I went to the GreenHab around 20:00 to the GreenHab with Aditya, our crew astronomer and Rajvi, our crew engineer. We performed the experiment namely ‘the pan experiment’ using some of the water from the GreenHab and left one pan in the GreenHab and kept one in the growth tent at science dorm. We will be checking them again today around the same time.
Sol 5 started a bit earlier than the other days. I woke up at 6 am and went to the GreenHab to water the plants. I did my usual duty of checking moisture levels for each crop and the crops were keeping up with the same levels so I decided to delay watering them to see if the moisture levels drops. The Crew woke up by 7:30 and started Yoga session. The whole crew woke up to stretch our bodies since some of us had to leave for EVA. Right after Yoga and breakfast I finished my mid mission report and left for EVA. Later in the evening around 18:30 I went to the hab and noticed the moisture level has dropped to 6-8%. So I watered all the crops and left back to hab.
Harvest: None
Support/supplies needed: None

EVA Report – February 10th

Crew 292 EVA Report 10-02-2024
EVA #007
Author: BHARTI SHARMA
Purpose of EVA: To collect the concretion sample, mark the locations of different types of weathering and understand the landscape at large for geo-conservation studies.
Start time: 13:00
End time: 16:00
Narrative: The weather cleared up today. However, there was snow on the ground but by the afternoon the weather was pleasant and the ground was not muddy at all. The EVA started at 13:00, on time. The crew walked towards the east direction from the Hab, in between the marble ritual and Pooh’s corner. The aim was to understand the landscape better for geo-conservation, collect concrete samples for biology and engineering projects, and identify and mark the locations for different types of weathering for geology projects. The crew collected some water samples as well for the biology project, using a syringe and falcon tubes. Clare was explaining in detail the geology of the area to Mehnaz, Aditya and Rajvi and they were quite intrigued. However, there was an issue with Rajvi’s suit, the fan couldn’t be lowered as the voltage dropped when they entered the airlock after finishing the EVA (refer to operations report). The crew was happy to be on the field, ticking off their field goals and collecting samples. As the crew was able to complete the EVA goals early, they came back to the Hab at 16:00.
Destination: Near the HAB, towards the east direction. It is between the Marble Ritual and Pooh’s corner, towards the northwest of the road. The field is within a radius of 700 m.
Coordinates (use UTM WGS 84): 518640, 4251224
Participants: Clare, Rajvi, Mehnaz and Aditya
Road(s) and routes per MDRS Map: Exited the HAB through the main airlock, and walked towards Cow Dung Road 0110 in the East direction. Walked further towards the east direction till the concretion point i.e., 518640, 4251224. Conducted the EVA within a radius of 700 m from the HAB.
Mode of travel: Walking

Supplemental Operations Report – February 10th

Date: 2/10/2024
Name of person filing report: Sergii Iakymov
Reason for Report: Routine
Non-Nominal Systems: HAB wall heater, solar observatory

Power system: Solar: Nominal.
SOC Last 24 hours: Max 100%; Min 76%; Avg 81.8%.
VDC Last 24 hours: Max 59.61V; Min 42.39; Avg 51.76V.
Generator run time: 5053.3 hours. Serviced on 02/08/2024, oil changed and air filter changed

Propane Readings: Station Tank: 66%
Director Tank: 60%
Intern Tank: 74%
Generator Tank: 30%

Water: Hab Static Tank – 341.09 gallons
GreenHab – 151.94 gallons
Outpost tank – 420 gallons
Science Dome – 0 gallons
Septic Tanks emptied: No

Rovers: Sojourner rover used: Yes, around campus
Hours: 196.6
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: Yes
Notes on Rovers: Perseverance is back to normal.

ATV: ATV’s Used: None. Nothing to report.

Cars: Hab Car used and why, where: To Hanksville for supplies. Need to be serviced.
Crew Car used and why, where: From Grand Junction by crew
General notes and comments: N/A.

Summary of Internet: All nominal.
EVA suits and radios: Suits: Suit 6 battery need to be tested. All others nominal
Comms: All nominal
Campus wide inspection, if action taken, what and why: All nominal.
Summary of Hab Operations: Wall heater is not functioning. Diagnostic is pending.
Summary of GreenHab Operations: All nominal.
Summary of SciDome Operations: All nominal.
Summary of Observatories Operations: Solar observatory telescope diagonal has two snapped screws. Repairs pending.
Summary of RAM Operations: All nominal.
Summary of Outpost Operations: Intern heater installed and running. Ventilation halls are made under the kitchen counter, bathroom sink and in the toilet door to help hot air move through. Rat holles have been patched,
Summary of Health and Safety Issues: All nominal.

Sol Summary – February 9th

Sol: Sol #5
Summary Title: Snow at MDRS
Author’s name: Annalea Beattie
Mission Status: Nominal
Sol Activity Summary: All field work is postponed today due to a layer of snow covering the desert. We regrouped after breakfast and met to set the goals for the day. Because of the weather forecast, Bharti and Clare spent some time planning the minimum schedule of EVAs necessary to complete everyone’s goals for field work in the following week. Bharti also read and took notes on the geomorphology of this area. She then prepared a base map. Aditya helped Mehnaz with evapotranspiration calculations for her research and then he had his scheduled shower. Later he continued his analysis of star images, specifically in response to the feedback given by Director of Observatories, Peter Detterline, and in relation to his image of NGC5904. Rajvi performed her engineer jobs and then discussed and examined variations of concretions with Clare. To minimize tools and save time, she is learning how to use her watch to mark points and trace her steps in the field. She also cooked a delicious lunch. Daniel spent much of the day in the Science Dome. In the morning he made two 5ml dilutions of the 5 different types of liquid media he brought with him. These were inoculated with 0.25g of the two different soil samples collected yesterday on EVA 4 and placed in the incubator which is set at 37C/99F to see if anything will grow. After lunch Daniel baked bread with Bharti and then extracted DNA from yesterday’s samples from EVA 4 in the Science Dome. I reorganized my notes for our upcoming outreach event with Women in Space, South Asia, did yoga and baked more bread. After two pm, I used the table and the magnification lamp in the Science dome to do several hours drawing and painting of concretion samples. Mehnaz gardened in the Green Hab, harvested greens for dinner and worked on the equations for her research.
Look Ahead Plan: Today while the snow melted we discussed our postponed EVAS and what might be possible for tomorrow, considering the forecast for snow, rain, thunderstorms and other hazardous weather predicted in the next few days. The forecast is not good for our fieldwork but we are trying to think about what is do-able and whether we can work close to the hab, weather permitting. Meanwhile we will ensure all our research is ready to go, as soon as the weather clears.
Anomalies in work: None
Weather: Snow covering the desert and sunny during the day with a high of 41F. Rain and cold with a wind chill factor of 30F and 22F overnight tonight with a chance of new snow accumulation of half an inch.
Crew Physical Status: Nominal. The crew are working well together as a team, everyone is healthy and morale is good.
EVA: None
Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA report, Astronomy Report, EVA request.
Support Requested: None

Operations Report – February 9th

SOL: #5
Name of person filing report: Rajvi Patel
Non-nominal systems: None
Notes on non-nominal systems: None
ROVERS
Spirit rover used: Not used
Opportunity rover used: Not used
Curiosity rover used: Not used
Perseverance rover used: Not used
General notes on rovers:
Summary of Hab operations: put notes here
WATER USE: 25.21 gallons (14.13 inches)
Water (static tank): 365.27 gallons
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: No
Summary of internet: Nominal
Summary of suits and radios: Nominal
Summary of GreenHab operations:
WATER USE: 7.031 gallons
Heater: On
Supplemental light: 22:00 – 02:00
Harvest: Radish (172g), Kale (96g), Red Cabbage (73g), Lettuce (33g)
Summary of Science Dome operations: In the morning, Daniel spent time in the Science Dome. He made two 5 ml dilutions of the 5 different types of liquid media he brought with him. These were inoculated with 0.25 g of the two different types of soil samples collected yesterday on EVA 4 and placed in the incubator which is set at 37 C 99F to see if anything will grow. Bharti used the Science Dome for her literature review on her research project. Annalea used the Science Dome to make drawings of concretions for her project.
Dual split: Heat On
Summary of RAM operations: (Tools used, work done) Went in to check if everything was okay due to the snowy weather.
Summary of any observatory issues: None
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support: None